How do we show that in the case of one dimensional elastic collision, the relative velocity of approach of 2 colliding bodies before collision is equal to the relative velocity of separation after collision?

Grade:11

1 Answers

Abishai M
7 years ago
By using energy equation (1/2)m1v1^2+(1/2)m2v2^2=(1/2)m1v1`^2+(1/2)m2v2`^2We get v1-v2=v2`-v1`Where v1 and v2 are the velocities of m1 and m2 respectivelyAlso for elastic collision coefficient of restitution is 1 hence we get the above separationV1` and v2` are velocities after collision of m1 and m2 respectively